(word problem) Jack and Joe leave their homes at the same time and drive toward each other. Jack drives at 60mph, while Joe drives at 30mph. They pass each other in 10 minutes. How far apart were Jack and Joe when they started?
Answer by mananth(16946)

Jack speed = 60mph
Joe speed = 30 mph
..
travelling towards each other
so speed = 30+60 =90 mph.
let distance be x when they start.
time take = 10 minutes = 1/6 hour
..
d= r*t
d= 90*1/6
d= 15 miles.
...
m.ananth@hotmail.ca